# A Strong Baseline for Temporal Video-Text Alignment

### v1

@@ Abstract
In this paper, we consider the problem of temporally aligning the video and texts from instructional videos, specifically, given a long-term video, and associated text sentences, our goal is to determine their corresponding timestamps in the video. To this end, we establish a simple, yet strong model that adopts a Transformer-based architecture with all texts as queries, iteratively attending to the visual features, to infer the optimal timestamp. We conduct thorough experiments to investigate: (i) the effect of upgrading ASR systems to reduce errors from speech recognition, (ii) the effect of various visual-textual backbones, ranging from CLIP to S3D, to the more recent InternVideo, (iii) the effect of transforming noisy ASR transcripts into descriptive steps by prompting a large language model (LLM), to summarize the core activities within the ASR transcript as a new training dataset. As a result, our proposed simple model demonstrates superior performance on both narration alignment and procedural step grounding tasks, surpassing existing state-of-the-art methods by a significant margin on three public benchmarks, namely, 9.3% on HT-Step, 3.4% on HTM-Align and 4.7% on CrossTask. We believe the proposed model and dataset with descriptive steps can be treated as a strong baseline for future research in temporal video-text alignment. All codes, models, and the resulting dataset will be publicly released to the research community.
1 Introduction
@@
The research on visual-language representation learning has recently made great progress. It is primarily driven by contrastive learning on image-caption pairs crawled from the Internet at scale [radford2021learning, li2022blip, li2023blip], and has shown remarkable performance on zero-shot image classification. However, in videos, the time dimension adds extra complexity that requires temporally corresponding captions/descriptions, posing challenges to learning fine-grained representation for video understanding tasks, such as temporal action localization, visual-language grounding, and grounded visual question answering.
@@
Instructional videos, for example, HowTo100M [miech2019howto100m], have been widely used for learning video representations [mavroudi2023learning, lin2022learning, han2022temporal, miech2020end], with the textual narrations acquired from Automatic Speech Recognition (ASR) system. However, unlike manually annotated captions, training model with these ASR transcripts naturally incurs three issues: (i) off-the-shelf ASR systems may introduce recognition errors; (ii) spoken narrations are generally not of descriptive style, thus contain redundancy or ambiguity, e.g., talking about a specific ingredient or using ambiguous pronouns (it, them); (iii) transcripts may not be well aligned with the visual signals, e.g., greetings from the speaker, associating with inaccurate timestamps, or describing the action after performing it. According to the statistics from the previous work [han2022temporal], only 30% of the narrations are visually alignable, and only 15% are naturally well-aligned with correct start/end timestamp, as presented in Fig. 1.
In this paper, we consider training a video-text alignment network on the large-scale instructional dataset – HowTo100M, where the associated ASR transcriptions have significant noise, and are only weakly aligned with the visual signal when relevant. Once trained, the model can tackle two challenging tasks, namely, narrations alignment [han2022temporal] and procedural steps grounding [mavroudi2023learning]. Specifically, given one long video sequence from the HowTo100M dataset, with associated narrations transcribed by the ASR system, or procedural steps from instructional articles on Wikihow111https://www.wikihow.com/Main-Page, that are related to the activity shown in the video, the narrations alignment task aims to: (i) justify if a sentence in narration is alignable with the video; (ii) determine its temporal segment if it is alignable. While the procedural steps grounding task aims to temporally ground the procedural steps in the video.
@@
We present a lightweight Transformer-based architecture, termed as Narrations / Steps to Video Aligner (NSVA), where the narration or procedural steps are used as queries, to iteratively attend the video features, and output the alignability or optimal temporal windows. Specifically, we start by investigating the effect of updating visual-textual backbones for video-text alignment, from S3D to CLIP, to the recent InternVideo. To resolve the aforementioned problems in ASR transcripts, we first replace the original YouTube transcripts with the ones generated from WhisperX [bain2022whisperx], to explore the effect of improving the ASR system in order to reduce the punctuation and recognition error in narrations, then we transform the ASR transcripts into descriptive steps, by prompting the large language model and temporally align them to the video through a two-stage determination procedure, i.e., approximate estimation, and then further refinement. As a result, we establish an automatic, scalable pipeline to construct high-quality data for training the video-text alignment network and obtain a large video-text dataset named HowToStep.
Our final model sets state-of-the-art performance on both procedural steps grounding and narrations alignment across three public benchmarks, surpassing existing models by a large margin, specifically, 9.3% on HT-Step [mavroudi2023learning], 3.4% on HTM-Align [han2022temporal] and 4.7% on CrossTask [zhukov2019cross]. We will release all codes, models, and the resulting dataset to facilitate future research.
@@ 2 Related Work
Large-Scale Video-Text Datasets. Multi-modal video datasets are crucial for video understanding tasks. Conventional video-text datasets [caba2015activitynet, zhou2018towards] are often manually labeled, suffer from short video lengths, limited scale, and coarse label granularity, and prevent the model from learning a generalized video representation. In the recent literature, scalability becomes an essential factor for constructing datasets, Youtube-8m [abu2016youtube] collects YouTube videos with metadata provided by the users, Instagram65M [ghadiyaram2019large] uses the associated hashtags as labels to supply weak supervision for training. In order to get descriptive sentences or captions with richer semantics, instructional videos are collected at scale, as they naturally come with dense narrations, obtained from ASR systems, by far, the largest video-text dataset is HowTo100M [miech2019howto100m]. As an alternative, Fit [Bain21] comprises over two million videos with weak captions scraped from the internet, while the captions are manually generated, they are not temporally aligned, and thus are insufficient for learning fine-grained temporal representation. In this work, we transform the noisy ASR transcripts into descriptive procedural steps, and propose to train a model to improve the video-text correspondence for instructional videos, mitigating the flaws of the HowTo100M dataset for visual representation learning.
@@
Video-Text Alignment. Video-text alignment aims to temporally assign narrations or procedural activity steps to the corresponding video segments. Initially, [bojanowski2014weakly, zhukov2019cross] tries to delineate the video segments corresponding to a given action list. Instead of an action list, scripts describing a series of events in the video are given for transcript alignment [cour2008movie, pramod2009subtitle]. The availability of large-scale video-text datasets such as HowTo100M has prompted many works on joint video-text embedding training. Specifically, TAN [han2022temporal] investigated directly aligning contextualized narration representations generated from ASR transcripts to video segments. Given that the ASR can be rather noisy, DistantSup [lin2022learning] proposes using distant supervision from a textual knowledge base, namely WikiHow [koupaee2018wikihow] to denoise text narrations from ASR. More recently, instead of aligning a video’s narrations, VINA [mavroudi2023learning] proposes to ground procedural activity steps sourced from the instructional article, however, as the order of instructions from Wikihow does not necessarily follow those in the video, grounding procedural steps is thus more challenging. In this paper, we aim to tackle the two problems simultaneously, namely, narration alignment [han2022temporal] and procedural step grounding [mavroudi2023learning], while previous studies have only focused on one aspect, leading to the inability to better integrate the two types of textual information (narrations and steps) to complete the video-text alignment tasks.
@@
Dataset Curation with Large Language Models. In the recent literature, large language models (LLMs) such as GPT [radford2019language, brown2020language, neelakantan2022text] and Alpaca [taori2023alpaca] have achieved great success in natural language processing. Constructing multi-modal datasets while generating pseudo-labels using LLMs becomes an efficient way to exploit the common sense knowledge in LLMs, and save human efforts for annotations. For instance, VQA-T [yang2021just] generates question-answer pairs for instructional video ASR transcripts by LLMs, which are then coupled with the related video. The VQA model pre-trained on the generated dataset exhibits enhanced generalization capabilities in various downstream tasks. Similarly, [changpinyo2022all] adopts LLMs to automatically derive question-answer pairs at scale for images, by leveraging the existing image-caption annotations. In addition to generating a dataset directly, some works use the LLMs to create pseudo-labels for large-scale video data that are later used for multi-modal vision tasks. For example, LAVILA [zhao2023learning] first trains a video captioning model on egocentric videos, and proposes to rephrase the densely generated captions from the model. After finetuning, the model demonstrates improved performance. As a concurrent work, [shvetsova2023howtocaption] adopts LLMs to transform the ASR transcripts of instructional videos, however, there exists one crucial difference, we transform the transcripts into concise steps, rather than dense captions as in their work.
@@ 3 Method
In this section, we start by describing the problem scenario in Sec. 3.1, including narration alignment and procedural step grounding. Then we present a novel pipeline to transform noisy and redundant ASR transcripts into descriptive steps as extra training source in Sec. 3.2. Lastly, we detail our proposed video-text alignment network in Sec. 3.3.
3.1 Problem Scenario
Given an instructional video from HowTo100M, , where refers to the frames of the video, and denotes the textual segments (narrations or procedural steps) associated with the video. Our goal is to train a temporal video-text alignment network that takes video and texts as inputs, and outputs a textual-visual alignment score matrix () with the binary visual alignability () for each sentence. Formally, it can be denoted as:
In training, the ground truth label takes value only if -th text depicts what is happening in timestamp in the video, and zero otherwise.
Depending on the type of given texts , we differentiate the two targeted tasks as follows:
Narration alignment [han2022temporal]. Narrations are textual sentences transcribed from the demonstrator’s speech via the ASR system, which means these sentences have maintained a time order that can be utilized for visual grounding. Note that, some narrations are not visually alignable or not well-aligned with visual signals temporally as shown in Fig. 1.
Procedural step grounding [mavroudi2023learning]. Procedural Steps are collected from Wikihow [koupaee2018wikihow], an external knowledge base with instructional articles explaining the procedures for completing certain tasks. Compared with narrations, the procedural steps do not necessarily have consistent temporal order as that happens in the video, or some steps may not even be present in the video at all. In addition, the training annotations for this task are scarce.
Discussion. The ASR transcript with timestamps can be used as weakly-aligned labels for training video-text alignment network as previous work [han2022temporal]. However, as narrations essentially aim to assist people with hearing impairment, thus are not required to be descriptive of visual scenes, and can be sometimes verbose with ambiguity. Training the alignment network to tackle both tasks only based on ASR transcripts is sub-optimal. While on the other hand, procedural steps are concise and clear, that only contain the core actions to finish one specific task without redundancy. Based on this observation, we propose a pipeline to transform the noisy ASR transcripts into concise and descriptive steps for instructional videos, by exploiting the power of large language models (LLMs), which effectively augment the dataset for training the alignment network, as detailed in the following subsection.
3.2 ASR Transcripts —¿ Aligned Descriptive Steps
@@
As presented in Fig. 2, the entire procedure can be divided into three parts: (i) we leverage LLM to summarize narrations from the ASR transcript into descriptive steps; (ii) we use the similarity between the original transcript and generated steps to roughly determine the start/end timestamp for each step as pseudo-label; (iii) we train the NSVA model on the generated steps with pseudo-label, and then use the trained model to refine the time range for each generated step (i.e., self-training). We name the final dataset as HowToStep and present the details in the following sections.
3.2.1 Prompting LLM to Summarize Concise Steps
@@
To generate descriptive steps, we split the complete transcript for each video into segments, each segment consists of around 10 sentences. Next, we prompt Llama2-7B [touvron2023llama] to summarize the transcripts into concise steps describing the main action presented in the video. Additionally, we prompt LLM to filter the colloquial sentences in the speech. We refer the readers to the complete prompt in Appendix.
@@
Formally, let refer to the large language model and the prompt texts, and the steps are generated separately:
@@
where is the number of separated ASR transcript segments. refers to the summarized procedure steps for each segment , which consists of zero to four descriptive steps. denotes the complete sequence of generated steps for one instructional video. As a result, we have transformed the ASR transcripts of around 370k videos (a subset of HowTo100M selected by [han2022temporal]) into approximately 7M descriptive steps. For comparison, the Wikihow knowledge base only contains 100k procedural steps from 14k instructional articles according to the statistics in [mavroudi2023learning].
3.2.2 Temporal Grounding for Generated Steps
As a baseline method, we can simply get the desired steps grounding matrix between steps and video timestamps by directly computing , where , is the corresponding visual-textual backbone (e.g., InternVideo). However, in practice, such an approach performs poorly, as shown in the ablation study at Tab. 4.
Alternatively, after generating descriptive steps, we propose to acquire their corresponding temporal windows in the instructional videos, with a two-stage approach to determine the start/end timestamp.
Stage 1: Approximate estimation. As the procedural steps are summarized from transcript sentences, we first compute the similarity score between generated steps and narrations of the transcript from each video, i.e., ‘stepstranscript’ matrix:
@@ 3.2.2 Temporal Grounding for Generated Steps
where is temperature, is the textual similarity matrix. The generated steps grounding matrix can then be computed by chaining the ‘stepstranscript’ and ‘transcriptvideo’ matrices, that has been defined in Sec. 3.1.
@@
We select the timestamp of the maximal alignment score as the center time for each step ,
@@
and then find start / end time () towards left and right side of until the alignment score () are lower than the percentage of the score of the center time (i.e., ) following [mavroudi2023learning]. The step with the maximal alignment score lower than the threshold will be regarded as unalignable and discarded.
Stage 2: Temporal refinement. In the first stage, the start / end time of the generated steps is completely obtained from weakly-aligned transcripts, which is naturally inaccurate, we further propose to refine the temporal windows with self-training. As shown in Fig. 2, we train the proposed video-text alignment network on the descriptive steps filtered after the first stage with their temporal windows as pseudo-labels. Then we use the trained model to do inference on the whole set of steps generated by LLM, to get new steps-to-video grounding matrix for each video, which might recycle some of the filtered steps back.
As briefly described in Sec. 3.1, when feeding video and a set of descriptive steps as input, the proposed NSVA model outputs the alignment matrix (). We take the timestamp of the maximal alignment score as the start time for each step, and the duration is a constant , following [chen2021multimodal, shvetsova2023howtocaption]. Similarly, the generated step with the maximal alignment score lower than the threshold will be discarded. Thus, we obtain a dataset consisting of aligned descriptive steps, named HowToStep. In practice, we observe that although our NSVA model is trained on noisy pseudo-labels (i.e., relying on weakly-aligned transcripts, and getting steps-transcript similarity from imperfect language model), the model tends to learn the alignment patterns from procedural step to videos before overfitting to the noises, as also being observed in [zhang2021understanding]. More details for generating the dataset are presented in ablation studies in Sec. 4.3.1 and Appendix.
Relation to existing work. Generally speaking, the proposed idea here is similar to the recent work [mavroudi2023learning], but with one crucial difference, that is, [mavroudi2023learning] aims to ground steps collected from the external knowledge base (i.e., Wikihow) as extra training source, while we try to align the descriptive steps summarized from the ASR transcript within the same video. This design brings two benefits for training the video-text alignment network: (i) the procedural steps generated per video are more diverse than those from Wikihow; (ii) the generated steps within the video are more likely to be alignable to the video itself than those externally sourced Wikihow steps.
3.3 Architecture

### v2

# https://lzq5.github.io/Video-Text-Alignment/ Multi-Sentence Grounding for Long-term Instructional Video

@@ Abstract
In this paper, we aim to establish an automatic, scalable pipeline for denoising the large-scale instructional dataset and construct a high-quality video-text dataset with multiple descriptive steps supervision, named HowToStep. We make the following contributions: (i) improving the quality of sentences in dataset by upgrading ASR systems to reduce errors from speech recognition and prompting a large language model to transform noisy ASR transcripts into descriptive steps; (ii) proposing a Transformer-based architecture with all texts as queries, iteratively attending to the visual features, to temporally align the generated steps to corresponding video segments. To measure the quality of our curated datasets, we train models for the task of multi-sentence grounding on it, i.e., given a long-form video, and associated multiple sentences, to determine their corresponding timestamps in the video simultaneously, as a result, the model shows superior performance on a series of multi-sentence grounding tasks, surpassing existing state-of-the-art methods by a significant margin on three public benchmarks, namely, 9.0% on HT-Step, 5.1% on HTM-Align and 1.9% on CrossTask. All codes, models, and the resulting dataset have been publicly released.
Keywords:
Instructional Video Understanding Video-Text Dataset Multi-Sentence Grounding1 Introduction
@@
The research on visual-language representation learning has recently made great progress. It is primarily driven by contrastive learning on image-caption pairs crawled from the Internet at scale [33, 21, 20], and has shown remarkable performance on zero-shot image classification. However, in videos, the time dimension adds extra complexity that requires temporally corresponding captions/descriptions, posing challenges to learning fine-grained representation for video understanding tasks, such as temporal action localization, visual-language grounding, and grounded visual question answering.
1) Descriptive: clearly describes the procedural action steps in the instructional video;
2) Concise: all sentences can be grounded in the video, without redundancy or noises;
3) Temporally well-aligned: offers precise temporal boundaries for procedural steps.
@@
Instructional videos, e.g., HowTo100M [29], have been widely used for learning video representations [27, 23, 17, 28], with the textual narrations acquired from the Automatic Speech Recognition (ASR) system. However, unlike manually annotated captions, training models with these ASR transcripts naturally incur three issues: (i) off-the-shelf ASR systems may introduce recognition errors; (ii) spoken narrations are generally not of descriptive style, thus contain redundancy or ambiguity, e.g., talking about a specific ingredient or using ambiguous pronouns (it, them, etc.); (iii) transcripts may not be well aligned with the visual signals, e.g., greetings from the speaker, associating with inaccurate timestamps, or describing the action before/after performing it. According to the statistics from [17], only 30% of the narrations are visually alignable, and only 15% are naturally well-aligned with correct start/end timestamp, as presented in Fig. 1.
@@
In this paper, we aim to establish an automatic, scalable pipeline for ‘denoising’ the large-scale instructional dataset, and contribute a ‘cleaned’ video-text dataset, termed as HowToStep. Specifically, our curation procedure involves the following aspects, first, to improve the quality of text descriptions, we replace the original YouTube transcripts with the ones generated from WhisperX [3], and prompt the large language models (LLMs) to transform noisy ASR transcripts into coherent, descriptive steps that are closely related to the video content; second, to temporally align the texts to the corresponding video timestamp, we adopt a two-stage determination procedure, i.e., approximate estimation based on timestamps from ASR transcripts, followed by training a lightweight Transformer-based model for further refinement, termed as Narrations / Steps to Video Aligner (NaSVA), where we use multiple sentences as queries, to iteratively attend the video features, and output the alignability or optimal temporal windows.
As a result, to demonstrate the effectiveness of our dataset curation procedure, we evaluate the model trained on such dataset, namely NaSVA, on multi-sentence grounding tasks - narrations alignment [17] and procedural steps grounding [27], both aiming to localize the corresponding temporal segments for multiple sentences in a video. Contrary to the existing text-to-video grounding tasks that focus on only a single sentence at a time, for example, temporal action localization [15], moment retrieval [19] and natural language queries [22], multi-sentence grounding necessitates the understanding of long instructional videos, with finer-grained events or actions, while managing multiple interrelated text queries simultaneously. Our model sets state-of-the-art performance on both procedural steps grounding and narrations alignment across three public benchmarks, surpassing existing models by a large margin, specifically, 9.0% on HT-Step [27], 5.1% on HTM-Align [17] and 1.9% on CrossTask [48].
@@ 2 Related Work
Large-Scale Video-Text Datasets. Multi-modal video datasets are crucial for video understanding tasks. Conventional video-text datasets [7, 47] are often manually labelled, suffer from short video lengths, limited scale, and coarse label granularity, and prevent the model from learning a generalized video representation. In the recent literature, scalability becomes an essential factor for constructing datasets, Youtube-8M [1] collects YouTube videos with metadata provided by the users, Instagram65M [16] uses the associated hashtags as labels to supply weak supervision for training. In order to get descriptive sentences with richer semantics, instructional videos are collected at scale, as they naturally come with dense narrations, obtained from ASR systems, by far, the largest video-text dataset is HowTo100M [29]. As an alternative, [4] comprises over two million videos with weak captions scraped from the internet, while the captions are manually generated, they are not temporally aligned, and thus are insufficient for learning fine-grained temporal representation. In this work, we transform the noisy ASR transcripts into descriptive procedural steps and propose to train a model to improve the video-text correspondence for instructional videos, mitigating the flaws of the HowTo100M dataset for visual representation learning.
@@
Language Grounding in Videos. Early efforts in language grounding predominantly concentrate on single sentence grounding tasks such as temporal action localization [15], moment retrieval [24, 45, 19], and natural language queries [22, 36], which lead to unprecedented progress in recent years. Given a temporally untrimmed video and a natural language query, the goal of single-sentence grounding is to determine the start and end times for the described content in the video. However, due to the limitation of query lengths to single sentences, the trained models lack an understanding of textual context. Unlike single-sentence grounding, multi-sentence grounding involves simultaneously grounding multiple sentences from different semantic scales within the video. The model needs to determine the intervals corresponding to each sentence based on the correlation between multiple sentences, significantly increasing the task’s complexity. This also enables the model to learn enhanced video-text representations. Initially, [5, 48, 14] tries to delineate the video segments corresponding to an action list. Instead of an action list, scripts describing a series of events in the video are given for transcript alignment [11, 32]. The availability of large-scale video-text datasets such as HowTo100M has prompted many works on joint video-text embedding training. Specifically, TAN [17] investigated directly aligning contextualized narration representations generated from ASR transcripts to video segments. Given that the ASR can be rather noisy, DistantSup [23] proposes using distant supervision from a textual knowledge base, namely Wikihow [18] to denoise narrations from ASR. More recently, instead of aligning narrations, VINA [27] proposes to ground procedural steps sourced from instructional article, however, as the order of instructions from Wikihow does not necessarily follow those in the video, grounding procedural steps is thus more challenging. In this paper, our improved HowToStep dataset enables to train models that tackle the two problems simultaneously, namely, narration alignment [17] and procedural step grounding [27], while previous studies have only focused on one aspect.
@@
Dataset Curation with Large Language Models. In the recent literature, large language models (LLMs) such as GPT [35, 6, 30] and Alpaca [38] have achieved great success in natural language processing. Constructing multi-modal datasets while generating pseudo-labels using LLMs becomes an efficient way to exploit the common sense knowledge in LLMs, and save human efforts for annotations. For instance, VQA-T [43] generates question-answer pairs for instructional video ASR transcripts by LLMs, which are then coupled with the related video. The VQA model pre-trained on the generated dataset exhibits enhanced generalization capabilities in various downstream tasks. Similarly, LLMs are adopted to automatically derive question-answer pairs at scale for images or videos, by leveraging the existing image/video captions [8, 12]. In addition to generating a dataset directly, some works use the LLMs to create pseudo-labels for large-scale video data that are later used for multi-modal vision tasks. For example, LAVILA [46] first trains a video captioning model on egocentric videos, and proposes to rephrase the densely generated captions from the model. After finetuning, the model demonstrates improved performance. As a concurrent work, [37] adopts LLMs to transform the ASR transcripts of instructional videos, however, there exists one crucial difference, we transform the transcripts into concise steps, rather than dense captions as in their work.
@@ 3 Method
In this section, our goal is to establish an automatic, scalable pipeline for ‘denoising’ the large-scale instructional dataset, and contribute a textually descriptive, and temporally-aligned video-text dataset, which is termed as HowToStep. As presented in Fig. 2, the entire pipeline can be divided into three parts: (i) we leverage LLM to transform narrations from the ASR transcript into descriptive procedural steps; (ii) we use the similarity between the original transcript and generated steps to approximate the start/end timestamp for each step; (iii) we train a lightweight multi-sentence grounding model on the generated steps with the approximated timestamp, and then use the trained model to refine the time range for each generated step (i.e., self-training).
In the following sections, we start with improving the quality of texts by LLMs in Sec. 3.1. Then, we introduce the multi-sentence grounding network that enables to align texts to their corresponding video segments in Sec. 3.2. Afterward, we describe the self-training procedure to refine the temporal ranges for the generated steps in Sec. 3.3. Lastly, we discuss two tasks that involve multi-sentence grounding, that can be addressed by our model, and reflect the quality of our curated dataset, namely, narration alignment and procedural step localization in Sec. 3.4.
3.1 Noisy ASR Transcripts Descriptive Procedural Steps
To improve the quality of texts of the ASR transcripts in HowTo100M, we first replace the original Youtube ASR transcripts with a recent WhisperX [3] as the speech recognition tool, which can potentially resolve errors in recognition or punctuation. However, some complex cases such as ambiguous pronouns and grammatical mistakes can not simply addressed by upgrading the ASR system. Therefore, we further propose to exploit the strong reasoning ability in large language models (LLMs), to summarize the procedural action steps from the noisy ASR transcripts, converting them to descriptive textual explanations for instructional videos. The details are as follows:
@@ 3.1 Noisy ASR Transcripts Descriptive Procedural Steps
Prompting LLM to Summarize Procedural Steps. We start by splitting the complete transcript for each video into segments, each segment consists of around 10 sentences. Next, we prompt a Llama2-7B [39] to summarize the transcript into concise steps that describe the main actions presented in the video, as well as filtering the colloquial sentences in the speech. We refer the readers to the complete prompt in Appendix.
@@
Formally, let refer to LLM and the prompt texts, and the steps are generated separately due to the limits of context length of LLM:
@@
where is the number of separated ASR transcript segments. refers to the summarized procedure steps for each segment . denotes the complete sequence of generated steps for one instructional video. As a result, we have transformed the ASR transcripts of around 370K videos (a subset of HowTo100M selected by [17]) into approximately 7M descriptive procedural steps. For comparison, the Wikihow knowledge base only contains 100k procedural steps from 14K instructional articles, according to the statistics in [27].
Estimating Timestamp For Procedural Steps. Till this point, we propose to equip the summarised procedural steps with corresponding video timestamps, by mapping them back to the original narrations with sentence similarity. Specifically, we compute the similarity score between generated steps and narrations of the transcript from each video, getting ‘stepstranscript’ matrix:
@@
where is a pre-trained text encoder, is temperature, and is the textual similarity matrix. The generated steps grounding score matrix can be computed by chaining the ‘stepstranscript’ matrix and ‘transcriptvideo’ matrix (i.e., ASR timestamps).
@@
We set the timestamp with maximal score as the centre time for each step ,
@@
and then find start-end timestamp () from until the alignment score () are lower than the percentage of the score of the center time (i.e., ) following [27]. The step whose maximal alignment score is lower than a threshold will be regarded as unalignable and discarded.
3.2 Multi-Sentence Grounding Model
At this point, the start-end timestamps of the generated steps are obtained directly from weakly-aligned transcripts, which are naturally inaccurate, here, we propose a multi-sentence grounding model, termed as Narrations / Steps to Video Aligner (NaSVA), and a self-training strategy to refine these timestamps. Specifically, given an untrimmed long-form instructional video (6.7 minutes on average), , where refers to the frames of the video, and denotes the textual sentences associated with the video. Our goal is to train a temporal multi-sentence grounding network that takes one video and multiple texts as inputs, and outputs a textual-visual alignment score matrix () with the binary visual alignability () for each sentence. Formally, it can be denoted as:
In training, the ground truth label takes value only if -th text depicts the scene of timestamp in the video, and zero otherwise.